A motorboat can go 10 miles downstream on a river in 20 min. Its takes 30 min.for this boat to go back upstream the in same 10 miles. Find the speed of the current.
Answer by ankor@dixie-net.com(15745)

A motorboat can go 10 miles downstream on a river in 20 min. Its takes 30 min.for this boat to go back upstream the in same 10 miles. Find the speed of the current.
:
Let x = speed of boat in still water
Let y = speed of the current
:
Change minutes to hrs: 20 min = .33 hrs; 30 min = .5 hrs
:
Find the actual speeds, Speed = Dist/time
Down stream: 10/.33 = 30 mph
Upstream: 10/.5 = 20 mph
:
Two equations:
x + y = 30
x - y = 20
------------adding eliminates y
2x + 0y = 50
x = 50/2
x = 25 mph is the boat speed in still water.
:
Find the current:
25 + y = 30
y = 5 mph is the current